OpenStudy (anonymous):

What is the solution set for |-4x - 8| = 4?

OpenStudy (mosaic):

First, simplify by dividing each term by 4. Then use the fact that if |a| = b, then a = +b or a = -b to solve for x.

OpenStudy (anonymous):

i understand the first part of what you said but the second part confuses me a bit @mosaic

OpenStudy (mosaic):

|-4x - 8| = 4 Divide throughout by 4 |-x - 2| = 1 |-x - 2| = 1 implies either -x-2 = +1 OR -x-2 = -1 Solve the two equations separately: -x-2 = 1 -x = 1+2 = 3 x = -3 OR -x-2 = -1 -x = -1+2 = 1 x = -1 x = -3 or -1.

OpenStudy (anonymous):

ohhhhhhhh ok thank you so much

OpenStudy (mosaic):

You are welcome.